OpenStudy (anonymous):

Find all solutions of the equation and express them in the form a + bi. x^2 − 5x + 7 = 0

OpenStudy (anonymous):

standard form is \[\frac{5}{2}+\frac{\sqrt{3}}{2}i\]and its conjugate \[\frac{5}{2}-\frac{\sqrt{3}}{2}i\]

OpenStudy (anonymous):

the a+bi for is also known as rectangular form. if you have a graphing calc you can type in your equation hit math go over to complex numbers and there is an option between polar and rectangular. pick rectangular.